# Structured Generative Modeling with The Thermodynamic Kolmogorov-Arnold Model

### v5

@@ 4 Conclusions
T-KAM provides a principled and scalable framework for structured and informed generative modeling. Our results demonstrate that correctly chosen basis functions and reference priors can enable more efficient training on simple or well-understood datasets through importance sampling. We also show how the Kolmogorov-Arnold representation theorem can inform architectural design and introduce structural bias, allowing T-KAM’s architecture to be guided entirely by the size of its latent space. Moreover, we find that the inverse transform method offers clear advantages over Langevin Monte Carlo sampling for energy-based priors in terms of speed and exact sampling. However, the benefits of annealing and population-based MCMC remain inconclusive, suggesting directions for further work. Future investigations into the practicality of our annealing methods depend on continued hardware advancements, such as those introduced by Zetta (2024).
@@ 4.1.1 Hardware acceleration
Additionally, while ITS enables fast inference and scales when the prior is a mixture model, T-KAM’s training with ULA does not scale efficiently on GPUs due to the iterative, gradient-based nature of ULA. GPUs are not well suited to Langevin Monte Carlo methods due to the irregular memory access patterns and bandwidth pressure caused by differentiating with respect to inputs, which requires storing the outputs of each layer within the generator.
Nonetheless, the unique advantages of T-KAM justify continued research and development. Moreover, Zetta (2024) introduces the Learnable Function Unit (LFU), a processor designed for efficient evaluation and differentiation of univariate, nonlinear functions in parallel, while offering precise data flow control that can mitigate the challenges of sparsity, recursion, and irregular memory access patterns. The LFU presents a promising avenue for more efficient implementations of ULA, and improved posterior sampling, potentially including bias correction and adaptive step size tuning with autoMALA, (Sec. A.6.4). Even without annealing, replacing ULA with autoMALA may yield a more robust probabilistic model with stronger theoretical guarantees.
Overall, T-KAM may prove more scalable than conventional generative models, owing to its reliance on finite representations of continuous functions and its compatibility with the LFU.
@@ 4.1.2 More opportunities for importance sampling
For SVHN and CelebA, IS proved infeasible as a posterior sampler, so ULA was used for Sec. 3.2. This may be attributed to prior–posterior mismatch, which rendered the IS estimator ineffective. Moreover, relying on IS restricts the prior to its initial domain, since proposals are generated from the prior using ITS and finite-domain KAN energy functions. In contrast, ULA can explore values of $\bm{z}$ beyond this domain, enabling adaptation of the prior domain within the grid-updating scheme for KANs proposed by Liu et al. (2024).
A possible remedy for the curse of dimensionality is to learn in a reduced-order space using Principal Component Analysis (PCA) (Jolliffe & Cadima, 2016), similar to the approach of Kingma et al. (2014), who also applied PCA to train VAEs. To address prior–posterior mismatch, future work might explore how existing domain knowledge can be incorporated to shape the prior’s energy functions and align them with the posterior before training begins. Additionally, efforts should be directed toward selecting appropriate initial domains or adapting the domains dynamically.
@@ 4.1.3 Expressivity
This might be achievable in T-KAM by introducing more latent distributions, or revising Eq. 13 into a Mixture of Experts framework (Cai et al., 2024) where $\alpha_{q,p}$ could be realized as a more flexible kernel to enable different inductive biases. However, this might be expensive for the model by Pang et al. (2020) since its use of Langevin sampling is computationally intense. For example, the previous work of Cui & Han (2024), which divides their latent EBM space using the diffusion strategy by Ho et al. (2020), states that sampling is expensive due to their use of MCMC methods.
T-KAM could potentially circumvent the additional cost by using IS as a fast sampler, though this would require addressing its current limitations.
@@ A.1 Codebase
The study was implemented at https://github.com/thezettascale/T-KAM.
@@ A.3.1 Cyclic annealing
The study by Calderhead & Girolami (2009) derives an analytic term for an optimal tempering distribution in the context of linear regression models, where the power posterior in Eq. 30 is used to evaluate the plausibility of model parameters in relation to the data, rather than reflecting a latent variable posterior. However, their findings may not hold in the context of T-KAM and generative modeling, making $p$ and the temperature schedule difficult variables to tune.
@@
Initializing $p\leq 1$ focuses learning on high-variance, data-informed likelihoods during the early stages of the cycle. Here, features emerge closer to the true posterior, and learning requires exploration and higher variance gradient estimates. As training progresses, T-KAM will have learned global features sufficiently well, reducing the meaningful change of measure between adjacent power posteriors of a $p\leq 1$ schedule. At this stage, $p(i)$ will have transitioned to $p>1$, thereby shifting focus towards exploitation and sculpting the prior with lower variance gradient estimates.
@@ A.4.7 Learning gradient for discretized Thermodynamic Integral and power posterior partition function
Telescoping the sum in Eq. 58, and noting that $k=1$ corresponds to the power posterior equaling the prior, and $k=N_{t}$ to the posterior, Eq. 58 yields T-KAM’s learning gradient:
@@ A.5 Hyperparameters and training details
spline_type (-) Cheby
spline_activation (-) None
@@
Initial basis param. $\tau_{0}$ 1.001
Train basis param. $\tau_{\text{trainable}}$ false
GaussQuad_nodes $N_{\text{quad}}$ 400
@@
Number of epochs $N_{\text{epochs}}$ 20
@@
iters $N_{\text{local}}$ 20
step_size $\eta$ 0.01
@@ A.5.1 SVHN & CelebA Architectures
1 $\times4$ convT($256$), LeakyReLU $4\times 4\times 256$ 1
2 $\times4$ convT($128$), LeakyReLU $8\times 8\times 128$ 2
3 $\times4$ convT($64$), LeakyReLU $16\times 16\times 64$ 2
@@ 4 $\times4$ convT(3), Sigmoid $32\times 32\times 3$ 2
1 $\times4$ convT($512$), LeakyReLU $4\times 4\times 512$ 1
2 $\times4$ convT($256$), LeakyReLU $8\times 8\times 256$ 2
3 $\times4$ convT($128$), LeakyReLU $16\times 16\times 128$ 2
4 $\times4$ convT(64), LeakyReLU $32\times 32\times 64$ 2
@@ A.6 Sampling theory
This section provides more theoretical detail regarding T-KAM’s sampling procedures.
@@ A.6.1 Importance Sampling
In the context of T-KAM, the target distribution is the posterior, $P(\bm{z}\mid\bm{x}^{(b)},\bm{f},\bm{\Phi})$, which is proportional to the product of the likelihood and the prior:
@@
The suitability of these weights depends on how well $Q(\bm{z})$ matches the posterior. In T-KAM, there is likely to be significant mismatch due to the complexity of the likelihood models described in Eq. 21. Consequently, the weights will exhibit high variance, resulting in many samples contributing minimally. This reduces the Effective Sample Size (ESS) and introduces bias into the estimates.
@@ A.7.4 SVHN
Figure 15: Generated SVHN examples by T-KAM trained with MLE and SE.
@@ A.7.5 CelebA
Figure 16: Generated CelebA examples by T-KAM trained with MLE and SE.
@@ A.8 Trained latent priors
The following priors were extracted from trained T-KAM models, each corresponding to the first three components of prior $q=1$.
@@ A.8.1 MNIST RBF
Figure 17: Three components from T-KAM’s prior after training on MNIST for 5,000 parameter updates with RBF bases and Gaussian constrained priors.
@@
Figure 18: Three components from T-KAM’s prior after training on MNIST for 5,000 parameter updates with RBF bases and lognormal constrained priors.
@@
Figure 19: Three components from T-KAM’s prior after training on MNIST for 5,000 parameter updates with RBF bases and uniform constrained priors.
@@ A.8.2 MNIST FFT
Figure 20: Three components from T-KAM’s prior after training on MNIST for 5,000 parameter updates with FFT bases and Gaussian constrained priors.
@@
Figure 21: Three components from T-KAM’s prior after training on MNIST for 5,000 parameter updates with FFT bases and lognormal constrained priors.
@@
Figure 22: Three components from T-KAM’s prior after training on MNIST for 5,000 parameter updates with FFT bases and uniform constrained priors.
@@ A.8.3 FMNIST RBF
Figure 23: Three components from T-KAM’s prior after training on FMNIST for 5,000 parameter updates with RBF bases and Gaussian constrained priors.
@@
Figure 24: Three components from T-KAM’s prior after training on FMNIST for 5,000 parameter updates with RBF bases and lognormal constrained priors.
@@
Figure 25: Three components from T-KAM’s prior after training on FMNIST for 5,000 parameter updates with RBF bases and uniform constrained priors.
@@ A.8.4 FMNIST FFT
Figure 26: Three components from T-KAM’s prior after training on FMNIST for 5,000 parameter updates with FFT bases and Gaussian constrained priors.
@@
Figure 27: Three components from T-KAM’s prior after training on FMNIST for 5,000 parameter updates with FFT bases and lognormal constrained priors.
@@
Figure 28: Three components from T-KAM’s prior after training on FMNIST for 5,000 parameter updates with FFT bases and uniform constrained priors.
@@ A.8.5 Darcy RBF
Figure 29: Three components from T-KAM’s prior after training on Darcy flow pressures for 10,000 parameter updates with RBF bases and Gaussian constrained priors.
@@
Figure 30: Three components from T-KAM’s prior after training on Darcy flow pressures for 10,000 parameter updates with RBF bases and lognormal constrained priors.
@@
Figure 31: Three components from T-KAM’s prior after training on Darcy flow pressures for 10,000 parameter updates with RBF bases and uniform constrained priors.
@@ A.8.6 Darcy FFT
Figure 32: Three components from T-KAM’s prior after training on Darcy flow pressures for 10,000 parameter updates with FFT bases and Gaussian constrained priors.
@@
Figure 33: Three components from T-KAM’s prior after training on Darcy flow pressures for 10,000 parameter updates with FFT bases and lognormal constrained priors.
@@
Figure 34: Three components from T-KAM’s prior after training on Darcy flow pressures for 10,000 parameter updates with FFT bases and uniform constrained priors.

### v6

@@ 4 Conclusions
KAEM provides a principled and scalable framework for structured and informed generative modeling. Our results demonstrate that correctly chosen basis functions and reference priors can enable more efficient training on simple or well-understood datasets through importance sampling. We also show how the Kolmogorov-Arnold representation theorem can inform architectural design and introduce structural bias, allowing KAEM’s architecture to be guided entirely by the size of its latent space. Moreover, we find that the inverse transform method offers clear advantages over Langevin Monte Carlo sampling for energy-based priors in terms of speed and exact sampling. However, the benefits of annealing and population-based MCMC remain inconclusive, suggesting directions for further work. Future investigations into the practicality of our annealing methods depend on continued hardware advancements, such as those introduced by Zetta (2024).
@@ 4.1.1 Hardware acceleration
Additionally, while ITS enables fast inference and scales when the prior is a mixture model, KAEM’s training with ULA does not scale efficiently on GPUs due to the iterative, gradient-based nature of ULA. GPUs are not well suited to Langevin Monte Carlo methods due to the irregular memory access patterns and bandwidth pressure caused by differentiating with respect to inputs, which requires storing the outputs of each layer within the generator.
Nonetheless, the unique advantages of KAEM justify continued research and development. Moreover, Zetta (2024) introduces the Learnable Function Unit (LFU), a processor designed for efficient evaluation and differentiation of univariate, nonlinear functions in parallel, while offering precise data flow control that can mitigate the challenges of sparsity, recursion, and irregular memory access patterns. The LFU presents a promising avenue for more efficient implementations of ULA, and improved posterior sampling, potentially including bias correction and adaptive step size tuning with autoMALA, (Sec. A.6.4). Even without annealing, replacing ULA with autoMALA may yield a more robust probabilistic model with stronger theoretical guarantees.
Overall, KAEM may prove more scalable than conventional generative models, owing to its reliance on finite representations of continuous functions and its compatibility with the LFU.
@@ 4.1.2 More opportunities for importance sampling
For SVHN and CelebA, IS proved infeasible as a posterior sampler, so ULA was used for Sec. 3.2. This may be attributed to prior-posterior mismatch, which rendered the IS estimator ineffective. Moreover, relying on IS restricts the prior to its initial domain, since proposals are generated from the prior using ITS and finite-domain KAN energy functions. In contrast, ULA can explore values of $\bm{z}$ beyond this domain, enabling adaptation of the prior domain within the grid-updating scheme for KANs proposed by Liu et al. (2024).
A possible remedy for the curse of dimensionality is to learn in a reduced-order space using Principal Component Analysis (PCA) (Jolliffe & Cadima, 2016), similar to the approach of Kingma et al. (2014), who also applied PCA to train VAEs. To address prior-posterior mismatch, future work might explore how existing domain knowledge can be incorporated to shape the prior’s energy functions and align them with the posterior before training begins. Additionally, efforts should be directed toward selecting appropriate initial domains or adapting the domains dynamically.
@@ 4.1.3 Expressivity
This might be achievable in KAEM by introducing more latent distributions, or revising Eq. 13 into a Mixture of Experts framework (Cai et al., 2024) where $\alpha_{q,p}$ could be realized as a more flexible kernel to enable different inductive biases. However, this might be expensive for the model by Pang et al. (2020) since its use of Langevin sampling is computationally intense. For example, the previous work of Cui & Han (2024), which divides their latent EBM space using the diffusion strategy by Ho et al. (2020), states that sampling is expensive due to their use of MCMC methods.
KAEM could potentially circumvent the additional cost by using IS as a fast sampler, though this would require addressing its current limitations.
@@ A.1 Codebase
The study was implemented at https://github.com/thezettascale/KAEM.
@@ A.3.1 Cyclic annealing
The study by Calderhead & Girolami (2009) derives an analytic term for an optimal tempering distribution in the context of linear regression models, where the power posterior in Eq. 30 is used to evaluate the plausibility of model parameters in relation to the data, rather than reflecting a latent variable posterior. However, their findings may not hold in the context of KAEM and generative modeling, making $p$ and the temperature schedule difficult variables to tune.
@@
Initializing $p\leq 1$ focuses learning on high-variance, data-informed likelihoods during the early stages of the cycle. Here, features emerge closer to the true posterior, and learning requires exploration and higher variance gradient estimates. As training progresses, KAEM will have learned global features sufficiently well, reducing the meaningful change of measure between adjacent power posteriors of a $p\leq 1$ schedule. At this stage, $p(i)$ will have transitioned to $p>1$, thereby shifting focus towards exploitation and sculpting the prior with lower variance gradient estimates.
@@ A.4.7 Learning gradient for discretized Thermodynamic Integral and power posterior partition function
Telescoping the sum in Eq. 58, and noting that $k=1$ corresponds to the power posterior equaling the prior, and $k=N_{t}$ to the posterior, Eq. 58 yields KAEM’s learning gradient:
@@ A.5 Hyperparameters and training details
spline_type (-) FFT
spline_activation (-) LeakyReLU
@@
Initial basis param. $\tau_{0}$ 1
Train basis param. $\tau_{\text{trainable}}$ true
GaussQuad_nodes $N_{\text{quad}}$ 200
@@
num_grid_updating_samples (-) 100
grid_update_frequency (-) 100
grid_update_decay (-) 0.9999
@@
Number of epochs $N_{\text{epochs}}$ 40
@@
iters $N_{\text{local}}$ 40
step_size $\eta$ 0.005
@@ A.5.1 SVHN & CelebA Architectures
1 $\times4$ convT($128$), LeakyReLU $4\times 4\times 128$ 1
2 $\times4$ convT($64$), LeakyReLU $8\times 8\times 64$ 2
3 $\times4$ convT($32$), LeakyReLU $16\times 16\times 32$ 2
@@ 4 $\times4$ convT(3), Sigmoid $32\times 32\times 3$ 2
1 $\times4$ convT($256$), LeakyReLU $4\times 4\times 256$ 1
2 $\times4$ convT($128$), LeakyReLU $8\times 8\times 128$ 2
3 $\times4$ convT($64$), LeakyReLU $16\times 16\times 64$ 2
4 $\times4$ convT(32), LeakyReLU $32\times 32\times 32$ 2
@@ A.6 Sampling theory
This section provides more theoretical detail regarding KAEM’s sampling procedures.
@@ A.6.1 Importance Sampling
In the context of KAEM, the target distribution is the posterior, $P(\bm{z}\mid\bm{x}^{(b)},\bm{f},\bm{\Phi})$, which is proportional to the product of the likelihood and the prior:
@@
The suitability of these weights depends on how well $Q(\bm{z})$ matches the posterior. In KAEM, there is likely to be significant mismatch due to the complexity of the likelihood models described in Eq. 21. Consequently, the weights will exhibit high variance, resulting in many samples contributing minimally. This reduces the Effective Sample Size (ESS) and introduces bias into the estimates.
@@ A.7.4 SVHN
Figure 15: Generated SVHN examples by KAEM trained with MLE and SE.
@@ A.7.5 CelebA
Figure 16: Generated CelebA examples by KAEM trained with MLE and SE.
@@ A.8 Trained latent priors
The following priors were extracted from trained KAEM models, each corresponding to the first three components of prior $q=1$.
@@ A.8.1 MNIST RBF
Figure 17: Three components from KAEM’s prior after training on MNIST for 5,000 parameter updates with RBF bases and Gaussian constrained priors.
@@
Figure 18: Three components from KAEM’s prior after training on MNIST for 5,000 parameter updates with RBF bases and lognormal constrained priors.
@@
Figure 19: Three components from KAEM’s prior after training on MNIST for 5,000 parameter updates with RBF bases and uniform constrained priors.
@@ A.8.2 MNIST FFT
Figure 20: Three components from KAEM’s prior after training on MNIST for 5,000 parameter updates with FFT bases and Gaussian constrained priors.
@@
Figure 21: Three components from KAEM’s prior after training on MNIST for 5,000 parameter updates with FFT bases and lognormal constrained priors.
@@
Figure 22: Three components from KAEM’s prior after training on MNIST for 5,000 parameter updates with FFT bases and uniform constrained priors.
@@ A.8.3 FMNIST RBF
Figure 23: Three components from KAEM’s prior after training on FMNIST for 5,000 parameter updates with RBF bases and Gaussian constrained priors.
@@
Figure 24: Three components from KAEM’s prior after training on FMNIST for 5,000 parameter updates with RBF bases and lognormal constrained priors.
@@
Figure 25: Three components from KAEM’s prior after training on FMNIST for 5,000 parameter updates with RBF bases and uniform constrained priors.
@@ A.8.4 FMNIST FFT
Figure 26: Three components from KAEM’s prior after training on FMNIST for 5,000 parameter updates with FFT bases and Gaussian constrained priors.
@@
Figure 27: Three components from KAEM’s prior after training on FMNIST for 5,000 parameter updates with FFT bases and lognormal constrained priors.
@@
Figure 28: Three components from KAEM’s prior after training on FMNIST for 5,000 parameter updates with FFT bases and uniform constrained priors.
@@ A.8.5 Darcy RBF
Figure 29: Three components from KAEM’s prior after training on Darcy flow pressures for 10,000 parameter updates with RBF bases and Gaussian constrained priors.
@@
Figure 30: Three components from KAEM’s prior after training on Darcy flow pressures for 10,000 parameter updates with RBF bases and lognormal constrained priors.
@@
Figure 31: Three components from KAEM’s prior after training on Darcy flow pressures for 10,000 parameter updates with RBF bases and uniform constrained priors.
@@ A.8.6 Darcy FFT
Figure 32: Three components from KAEM’s prior after training on Darcy flow pressures for 10,000 parameter updates with FFT bases and Gaussian constrained priors.
@@
Figure 33: Three components from KAEM’s prior after training on Darcy flow pressures for 10,000 parameter updates with FFT bases and lognormal constrained priors.
@@
Figure 34: Three components from KAEM’s prior after training on Darcy flow pressures for 10,000 parameter updates with FFT bases and uniform constrained priors.